OpenStudy (anonymous):

Find the measure of an angle whose measure is 18 degrees less than one-half of the measure of its complement. If x represents the measure of the angle, then which of the following expressions would equal x? 90 - (0.5x - 18) 0.5(90 - x - 18) 0.5(90 - x) - 18

OpenStudy (***[isuru]***):

if y and x r complementary angle x + y = 90 now the question says the angle we need is 18 degree less than the one half of the measure of its complement.. if the angle we need to find is x then its compliment is "y" so x = 0.5 ( y - 18 ) but we know x + y = 90 so.. y= 90 - x which means x = 0.5 ( 90 - x - 18) hope this will help ya!! :)